OpenStudy (anonymous):

solve y = (x+8)(x+4)(x-6)

OpenStudy (anonymous):

when solving for 'x', take each factor, set it equal to zero, then solve for x. i'll do the first so you get the idea: (x+8) (x+8)=0 x=-8 ur turn, do the same steps for the other two factorized terms: (x+4) and (x-6)

OpenStudy (anonymous):

(+4) (x+4)=0 x=-4 @DemolisionWolf

OpenStudy (anonymous):

correct, ^_^ now just do the last one, (x-6)

OpenStudy (anonymous):

(+6) (+6)=0 x=-6 @DemolisionWolf

OpenStudy (anonymous):

correct, just don't forget to write the 'x's in, like this: (x+6) (x+6)=0 x=-6

OpenStudy (anonymous):

can i tag you in something and maybe you could help?

OpenStudy (anonymous):

sure thing